OpenStudy (anonymous):

Solve the following equation. 5(x - 4) = 12 x=

OpenStudy (anonymous):

5(x-4) = 12 5x - 20 = 12 5x = 34 x = 34/5

OpenStudy (anonymous):

Sorry, 5x = 32 x = 32/5